OpenStudy (anonymous):

simplify: 2(-11x-7) (a) -22x - 14 (b) -11x-14 (c)-22-14x (d) 22x-14

OpenStudy (anonymous):

a negative times a postive is a negative so multiply: 2x-11= ? 2x-7=?

OpenStudy (anonymous):

idk why it says im typing bc im not

OpenStudy (anonymous):

2x-7=11

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no whats 2 times 7?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

7x1=7 7x2=14 so since you have -7 it would be -14

OpenStudy (anonymous):

lol they all have -14

OpenStudy (anonymous):

basically with 2(-11x-7), you are going to multiply both of the numers by 2